OpenStudy (anonymous):

The slope of a line is 8 and the point (-6, -16) is on the line. Express the line in slope-intercept form. Select the best answer from the choices provided. y = -16x + -104 y = 8x + 32 y = 8x + 122 y = -16x + 32

OpenStudy (texaschic101):

slope intercept form : y = mx + b .... m is your slope and b is your y intercept y = mx + b slope(m) = 8 (-6,-16)....x = -6 and y = -16 we already know slope(m) is 8, so now we need to find the y intercept (b) now we sub -16 = 8(-6) + b -16 = -48 + b -16 + 48 = b 32 = b so your equation is : y = 8x + 32 any questions ?